# Prevention is better than cure? Feedback from high specific energy winds in cosmological simulations with Arkenstone

### v1

@@ 3.5 Evolution with redshift
To emphasise the differences in redshift evolution between the runs we plot the global star formation rate density (SFRD) versus redshift in Fig. 14. It should be noted that comparison to observations (included for reference) should be made with caution because of the limited box size. However, the relative comparison between simulations remains informative. In these panels, it is clear that cosmic star formation history is most sensitive to the change in energy loading.
In the top panel, with fixed mass loadings of $\eta_{\rm M}=0.3$, the run with the lowest energy loading, $\eta_{\mathrm{E}}=0.1$ has a global SFRD that rises higher than any other run at cosmic noon. And while the higher energy loading runs generally have a declining SFRD to $z=0$, the lowest energy loading simulation turns up again at low redshift. The Arkenstone simulations all tend to have the turnover at cosmic noon occurring slightly earlier than in TNG, though we caution this is dominated by the largest few haloes in the box. These are in turn sensitive to the interaction between stellar feedback and the TNG BH model, which again has not been calibrated and is beyond the scope of this work.
For runs with a fixed energy loading $\eta_{\rm E}=0.3$, varying the mass loading of the hot winds also affects the SFRD, though for high and intermediate redshifts the effect is smaller than that of energy loading. Again, lower mass loadings suppress star formation further, due to the wind being hotter and faster (e.g Voit et al., 2024a). For the case of $\eta_{\mathrm{M}}=0.9$, the higher mass loading of the wind has disastrous consequences for the regulation of star formation, with the star formation rate rising essentially unchecked after $z=0.3$. We stop this simulation shortly after this, as the high densities of gas, stars and wind particles lead to a significant increase in computational cost.
The SFRDs for runs with variable energy loading have a different shape to that of TNG, with a suppression of star formation at high redshift, before a rapid increase to cosmic noon. This is due to the higher energy loading for lower halo masses, which suppress star formation at early times. As discussed in Section 4, the tests we have performed in this work have not been calibrated, and are an investigation of how high specific energy winds alone can regulate star formation via preventative feedback. To fully explain the presence of the massive early galaxies identified with JWST (e.g. Finkelstein et al., 2024) may require a change in the star formation prescription and possibly a combination of ejective and preventative feedback, which we will study in detail in future work with the full Arkenstone model.
Figure 14: Global star formation rate density at each simulation timestep in our simulated boxes as a function of redshift. Line styles are the same as those in Fig. 5. Observationally inferred data comes from several different methods (Robertson & Ellis, 2012; Fudamoto et al., 2021; Gruppioni et al., 2020; Cochrane et al., 2023; Algera et al., 2023). Higher energy loadings can lead to much more suppression in star formation.
@@ 4 Discussion
The core result of this paper is that star formation can be regulated by high specific energy winds. Fig. 4 shows this across a wide variety of halo masses using such winds only (as a reminder, the cold cloud component of the full Arkenstone framework is not active in this paper). This is a significant pivot away from the highly mass loaded winds of many existing cosmological simulations – a shift from primarily ejective towards primarily preventative feedback. The balance between the two has yet to be determined, and will require cosmological simulations with the full Arkenstone model, but we have shown in this work how high mass loadings are not necessary to regulate star formation in low and intermediate mass galaxies if sufficient energy is injected into the system via hot winds (controlled by the ratio of energy loading to mass loading). This does not require unphysically large amounts of energy to be injected, indeed we emphasise that Arkenstone is green: star formation is regulated while using considerably less energy than existing simulations (see Fig. 1). The important characteristic demonstrated in this work is the high specific energy of the wind, further discussed below (and in Voit et al., 2024a, b). Second-order observables, especially of the CGM, will be essential to disentangle the contribution from ejective and preventative feedback in the future.
As discussed in Section 2.6, at $z=0$ the mass loadings of galaxies at $10^{10}$, $10^{11}$, and $10^{12}$ M in the TNG model are approximately 31.4, 10.2 and 1.5, respectively. For energy loadings the values are approximately 3.6, 1.36 and 0.93, respectively. Compare these to the fixed $\eta_{\mathrm{M}}=\eta_{\mathrm{E}}=0.3$ of the central fixed variation of Arkenstone  or even the variable energy loading run which has a cap at $\eta_{\mathrm{E,max}}=1$, and we can see a dramatic shift in how galaxies are regulated – much less material enters and leaves the galaxy with Arkenstone’s preventative feedback.
@@
In Sections 1 and 2 we motivated our investigation and choices of parameters based on recent analytic and semi-analytic works. The gas-regulator model of Carr et al. (2023) simplifies the baryon cycle to an exchange of mass, metals, and energy between reservoirs of stars, the ISM and the CGM. They find that observables such as the SMHM relation can be reproduced with winds that have low mass loadings and high specific energy. They find that $z=0$ stellar masses are largely insensitive to changes in mass (and metal) loading, but dependent on energy loading, in good agreement with our results. While the dependence of $\eta_{\rm E}$ on halo mass that Carr et al. (2023) finds is needed to reproduce the SMHM is slightly steeper than what we have tested, we find the same general trend of higher $\eta_{\rm E}$ at lower halo masses is required. A notable result of Carr et al. (2023) is that highly mass loaded winds tend to lead to a denser CGM with lower specific energy, allowing gas to return to the galaxy on shorter timescales. Our results on the CGM gas fraction and temperature support this, with the high mass loaded winds of TNG generally leading to a cooler, more massive CGM (see Figs. 8, 9, 12, and 13).
@@
Pandya et al. (2023) used a semi-analytic model framework to study CGM evolution and its impact on galaxy evolution, with the additional inclusion of turbulence to act as regulation in addition to thermal energy. Like us, Pandya et al. (2023) finds that the average global CGM temperature does not have to be at the virial temperature (and is in fact often below). They also find higher specific energy winds (with higher $\eta_{\rm E}/\eta_{\rm M}$) to lead to a hotter, more turbulent CGM. We do find a similar temperature increase at $z=2$ (Fig. 13), though the trend is less clear at $z=0$. Interestingly, with Arkenstone the ratio $\eta_{\rm E}/\eta_{\rm M}$ alone does not set the average CGM temperature, as the runs with ($\eta_{\rm E}$, $\eta_{\rm M}$) = (0.9, 0.3) and (0.3, 0.1) tend to have slightly different temperatures (with the former being hotter) despite the same loading ratio. We postpone a detailed study of CGM turbulence to future work.
@@
A key comparison to make is to the minimalist regulator model of Voit et al. (2024a, b), which reduces galaxy evolution and the baryon cycle to three coupled differential equations. Arkenstone acts to couple supernova feedback to the CGM. This model predicts that the CGM expands and contracts depending on the comparison between the specific energy of gas injected into the CGM via feedback versus the specific energy it had when accreted. Thus, the minimalist regulator model offers a physical interpretation for many of the results in this paper – high specific energy winds drive the expansion of the CGM, reducing the inflow rate onto central galaxies and the corresponding SFR. This is particularly evident in Fig. 3, which highlights the expanded size of the CGM around many galaxies in our simulated volume. The regulator model also offers an explanation for why our Arkenstone simulation with the highest mass loading catastrophically fails at low redshift - a higher $\eta_{\rm M}$ at fixed $\eta_{\rm E}$ drops the specific energy of the wind, leading to a contracted CGM, more cooling, more recycling of material, and thus a higher SFR. Our simulations provide evidence for the important role of the specific energy of galactic winds in regulating star formation and shaping the properties of the CGM.
@@ 5 Conclusions
In this paper we have demonstrated the use of the new Arkenstone galactic wind model in cosmological simulations for the first time. We described the refinement scheme required for the Arkenstone-Hot model to work in cosmological simulations, before testing it on a (36.936.936.9\,36.9Mpc)3 box. We perform a series of numerical experiments, with both fixed mass and energy loadings and variable energy loadings that scale with DM velocity dispersion (a proxy for halo mass). In this work we do not calibrate our results to observations, we simply explore the impact of changing the loadings. Throughout the paper, we present our results next to the TNG model for reference.
Our key findings are as follows:
@@
    The value of energy loading is the key parameter which affects the stellar masses and sizes of galaxies and properties of the CGM, with a much smaller change arising from variations in mass loading – both the total energy injected into the CGM and the specific energy of the wind contribute to these results.
@@
    Arkenstone can match observed galaxy scaling relations without exceeding the supernovae energy budget or the low mass loadings inferred from observations of dwarf galaxy outflows.
Future work with Arkenstone will delve further into the new parameter space opened up by the inclusion of detailed modelling of high specific energy winds. An important aspect of the Arkenstone-Hot model to be refined further is the metal loading of the hot winds, which will have a significant impact on observables of the CGM and IGM. Additionally, while in this work we tied wind loadings to halo–scale properties (enabling a simple numerical experiment), in future we will instead base them on local ISM properties (informed by small scale ISM patch simulations), increasing predictive power. Another key advance to be made with Arkenstone is the utilisation of the full multiphase wind model (Smith et al., 2024b) in cosmological volumes, with subgrid cool clouds embedded in the hot wind. Only by using both aspects of the model can we begin to determine the balance between ejective and preventative feedback in galaxy evolution.
@@ Acknowledgements
JSB wishes to thank Laura Sommovigo and rachel somerville for assembling observational comparison data, and Stephen Thorp and Justin Alsing for sharing data from pop-cosmos (Alsing et al., 2024). The authors would also like to thank Shy Genel for providing the initial conditions used in these simulations, and Chris Carr and Mark Voit for useful conversations. This work was supported by the Simons Collaboration on “Learning the Universe”. Computations were performed on the HPC systems Iron, at the Flatiron Institute, Orion, at the Max Planck Computing and Data Facility (MPCDF), and Anvil, at Purdue University. The latter was accessed through allocation PHYS240043 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296 (Song et al., 2022; Boerner et al., 2023). GLB acknowledges support from the NSF (AST-2108470, AST-2307419), NASA TCAN award 80NSSC21K1053, the Simons Foundation. This work made use of the NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), and CMasher (van der Velden, 2020) Python packages.

### v2

@@ 3.5 Evolution with redshift
To emphasise the differences in redshift evolution between the runs we plot the global star formation rate density (SFRD) versus redshift in Fig. 16. It should be noted that a comparison to observations is not trivial to make, not least because of the limited box size, so we omit this here. However, the relative comparison between simulations remains informative. In these panels, it is clear that cosmic star formation history is most sensitive to the change in energy loading.
In the top panel, with fixed mass loadings of $\eta_{\rm M}=0.3$, the run with the lowest energy loading, $\eta_{\mathrm{E}}=0.1$ has a global SFRD that rises higher than any other run at cosmic noon, and its turnover occurs slightly earlier than in TNG, though we caution this is dominated by the largest few haloes in the box. These are in turn sensitive to the interaction between stellar feedback and the TNG BH model, which again has not been re-calibrated and is beyond the scope of this work. The highest energy loading has a significantly lower SFR throughout cosmic history - especially at high redshift and down to cosmic noon.
For runs with a fixed energy loading $\eta_{\rm E}=0.3$, varying the mass loading of the hot winds also affects the SFRD, though for high and intermediate redshifts the effect is smaller than that of energy loading. Again, lower mass loadings suppress star formation further, due to the wind being hotter and faster (e.g Voit et al., 2024a).
The SFRDs for runs with variable energy loading have a different shape to those from TNG, with a suppression of star formation at high redshift, before a rapid increase to cosmic noon. This is due to the higher energy loading for lower halo masses, which suppress star formation at early times. As discussed in Section 4, the tests we have performed in this work have not been calibrated, and are an investigation of how high specific energy winds alone can regulate star formation via preventative feedback. To fully explain the number densities of the UV luminous early galaxies identified with JWST (e.g. Bunker et al., 2023; Curtis-Lake et al., 2023; Tacchella et al., 2023; Finkelstein et al., 2024) may require a change in the star formation prescription and possibly a combination of ejective and preventative feedback, which we will study in detail in future work with the full Arkenstone model.
Figure 16: Global star formation rate density at each simulation timestep in our simulated boxes as a function of redshift. Line styles are the same as those in Fig. 5. It is not trivial to make observational comparisons to this quantity, so we postpone this to future work. Higher energy loadings can lead to much more suppression in star formation.
@@ 4 Discussion
The core result of this paper is that star formation can be regulated by high specific energy winds, with low input mass loadings and sub-unity energy loadings. Fig. 4 shows this across a wide variety of halo masses using such winds only (as a reminder, the cold cloud component of the full Arkenstone framework is not active in this paper). This is a significant pivot away from the highly mass loaded winds of many existing cosmological simulations – a shift from primarily ejective towards primarily preventative feedback. The balance between the two has yet to be determined, and will require cosmological simulations with the full Arkenstone model, but we have shown in this work how high mass loadings are not necessary to regulate star formation in low and intermediate mass galaxies if sufficient energy is injected into the system via hot winds (controlled by the ratio of energy loading to mass loading). Further, less easily accessibly observables, especially of the CGM, will be essential to disentangle the contribution from ejective and preventative feedback in the future.
Arkenstone has low mass loadings, but importantly also does not require unphysically large amounts of energy to be injected. Indeed we emphasise that Arkenstone is green: star formation is regulated while using considerably less energy than existing simulations. As discussed in Section 2.6, at $z=0$ the mass loadings of galaxies at $10^{10}$, $10^{11}$, and $10^{12}$ M in the TNG model are approximately 31.4, 10.2 and 1.5, respectively. For energy loadings the values in TNG in haloes of $10^{10}$, $10^{11}$, and $10^{12}$ M are approximately 3.6, 1.36 and 0.93, respectively. Compare these to the fixed $\eta_{\mathrm{M}}=\eta_{\mathrm{E}}=0.3$ of the central fixed variation of Arkenstone  or even the variable energy loading run which has a cap at $\eta_{\mathrm{E,max}}=1$, and we can see a dramatic shift in how galaxies are regulated – much less material enters and leaves the galaxy with Arkenstone’s preventative feedback. We also note that EAGLE requires an enhanced energy injection (likely somewhat higher than what is nominally available from supernovae) to regulate low mass galaxies (see e.g. Crain et al., 2015). As discussed in Section 2.6, the actual amount of energy available to drive winds out of the galaxy will be smaller than $\eta_{\mathrm{E,max}}$, as energy will be needed to escape the ISM. Energy loadings larger than unity are likely to be unphysical, but the true amount of energy available to drive winds must be investigated in small-scale simulations.
@@
In Sections 1 and 2 we motivated our investigation and choices of parameters based on recent analytic and semi-analytic works. The gas-regulator model of Carr et al. (2023) simplifies the baryon cycle to an exchange of mass, metals, and energy between reservoirs of stars, the ISM and the CGM. They find that observables such as the SMHM relation can be reproduced with winds that have low mass loadings and high specific energy. They find that $z=0$ stellar masses are largely insensitive to changes in mass (and metal) loading, but dependent on energy loading, in good agreement with our results. While the dependence of $\eta_{\rm E}$ on halo mass that Carr et al. (2023) find is needed to reproduce the SMHM is slightly steeper than what we have tested, we find the same general trend of higher $\eta_{\rm E}$ at lower halo masses is required. A notable result of Carr et al. (2023) is that highly mass loaded winds tend to lead to a denser CGM with lower specific energy, allowing gas to return to the galaxy on shorter timescales. Our results on the CGM gas fraction and temperature support this, with the highly mass loaded winds of TNG generally leading to a cooler, more massive CGM (see Figs. 8, 10, 14, and 15).
@@
Pandya et al. (2023) used a semi-analytic model framework to study CGM evolution and its impact on galaxy evolution, with the additional inclusion of turbulence in addition to thermal energy. Consistent with the work presented here, Pandya et al. (2023) find that the average global CGM temperature does not have to be at the virial temperature (and is in fact often below). They also find higher specific energy winds (with higher $\eta_{\rm E}/\eta_{\rm M}$) to lead to a hotter, more turbulent CGM. We do find a similar temperature increase at $z=2$ (Fig. 15), though the trend is less clear at $z=0$. Interestingly, with Arkenstone the ratio $\eta_{\rm E}/\eta_{\rm M}$ alone does not set the average CGM temperature, as the runs with ($\eta_{\rm E}$, $\eta_{\rm M}$) = (0.9, 0.3) and (0.3, 0.1) tend to have slightly different temperatures (with the former being hotter) despite the same loading ratio. We postpone a detailed study of the role of CGM turbulence to future work.
@@
A key comparison is to the minimalist regulator model of Voit et al. (2024a, b), which reduces galaxy evolution and the baryon cycle to three coupled differential equations. This model predicts that the CGM expands and contracts depending on the ratio between the specific energy of gas injected into the CGM via feedback versus the specific energy it had when accreted. Thus, the minimalist regulator model offers a physical interpretation for many of the results in this paper – high specific energy winds drive the expansion of the CGM, reducing the inflow rate onto central galaxies and the corresponding SFR. This is particularly evident in Fig. 3, which highlights the expanded size of the CGM around many galaxies in our simulated volume. The regulator model also offers an explanation for why our Arkenstone simulation with the highest mass loading fails to regulate star formation - a higher $\eta_{\rm M}$ at fixed $\eta_{\rm E}$ reduces the specific energy of the wind, leading to a contracted CGM, more cooling, more recycling of material, and thus a higher SFR. Our simulations provide evidence for the important role of the specific energy of galactic winds in regulating star formation and shaping the properties of the CGM.
@@ 5 Conclusions
In this paper we have demonstrated the use of the new Arkenstone galactic wind model in cosmological simulations for the first time. We described the refinement scheme required for the Arkenstone-Hot model to work in cosmological simulations, before testing it on a (36.936.9\,Mpc)3 box. We perform a series of numerical experiments, with both fixed mass and energy loadings and variable energy loadings that scale with DM velocity dispersion (a proxy for halo mass). In this work we do not calibrate our results to observations, but simply explore the impact of changing the loadings. Throughout the paper, we present our results in comparison with simulations run with the TNG model for reference.
Our key conclusions are as follows:
@@
    The energy loading is the key parameter which affects the stellar masses and sizes of galaxies and properties of the CGM, with a much smaller change arising from variations in mass loading – both the total energy injected into the CGM and the specific energy of the wind contribute to these results.
@@
    Arkenstone can match observed galaxy scaling relations without exceeding the supernovae energy budget and with significantly lower mass loadings than required in other cosmological simulations.
Future work with Arkenstone will delve further into the new parameter space opened up by the inclusion of detailed modelling of high specific energy winds. An important aspect of the Arkenstone-Hot model to be refined further is the metal loading of the hot winds (discussed in Appendix A), which will have a significant impact on observables of the CGM and IGM. Additionally, while in this work we tied wind loadings to approximately halo–scale properties (enabling a simple numerical experiment), in future we will instead base them on local ISM properties (informed by small scale ISM patch simulations), increasing predictive power. Another key advance to be made with Arkenstone is the utilisation of the full multiphase wind model (Smith et al., 2024b) in cosmological volumes, with subgrid cool clouds embedded in the hot wind. Only by using both aspects of the model can we begin to determine the balance between ejective and preventative feedback in galaxy evolution.
@@ Acknowledgements
The authors thank the anonymous referee for helpful comments that have led to an improved paper. The authors would like to thank Shy Genel for providing the initial conditions used in these simulations, and Chris Carr and Mark Voit for useful conversations. This work was supported by the Simons Collaboration on “Learning the Universe”. The authors wish to thank the synthetic observations working group of the Learning the Universe collaboration for assembling observational comparison data, and Stephen Thorp and Justin Alsing for sharing data from pop-cosmos (Alsing et al., 2024). Computations were performed on the HPC systems Iron, at the Flatiron Institute; Orion, at the Max Planck Computing and Data Facility (MPCDF); Anvil, at Purdue University (accessed through allocation PHYS240043 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296); and Frontera (accessed through Leadership Resource Allocation AST20007), at the Texas Advanced Computing Center (Stanzione et al., 2020; Song et al., 2022; Boerner et al., 2023). GLB acknowledges support from the NSF (AST-2108470, AST-2307419), NASA TCAN award 80NSSC21K1053, and the Simons Foundation. The Flatiron Institute is supported by the Simons Foundation. This work made use of the numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), and cmasher (van der Velden, 2020) python packages.
@@
Appendix A Metallicity of galactic winds
Figure 17: Metallicity of stars within twice the stellar half mass radius, as a function of stellar mass. We show TNG (black) and our central fixed loading Arkenstone run ($\eta_{\mathrm{E}}=\eta_{\mathrm{M}}=0.3$ (orange), and additionally show the same Arkenstone run but with a the wind metallicity enhanced by a factor of 5 (red).
Figure 18: Median gas metallicity (within twice the stellar half mass radius) as a function of stellar mass, with the line styles the same as Fig. 17. For comparison we show observational data from Tremonti et al. (2004), with the dotted line showing the median, and dark and light grey shaded regions encompassing the 1- and 2-$\sigma$ scatter, respectively.
Figure 19: Stellar mass - halo mass relation for TNG, Arkenstone with fixed $\eta_{\mathrm{E}}=\eta_{\mathrm{M}}=0.3$, and the same Arkenstone run with wind metallicity enhanced by a factor of 5. The stellar mass of galaxies does not change significantly with variations in wind metallicity, especially compared with variations due to loading parameters.
The metal content of galaxies is an important diagnostic for the amount of material produced by stars that remains within the galaxy as it evolves, as metals created in a galaxy can only be locked into future stars or leave via outflows. Stellar metallicity provides an integrated version of this, with enriched gas remaining in a galaxy forming an enriched generation of stars. Gas metallicity is a more variable measurement, and can be affected by both enriched outflows and fresh inflows from the IGM.
The mass loading of winds directly affects the metallicity of a galaxy, as it determines how much gas is removed from the ISM. The very high mass loadings of TNG in the dwarf regime therefore have to be offset by ‘wind stripping’, whereby metals are removed from wind particles at launch and distributed over neighbouring ISM cells, to decrease the metallicity of the wind and avoid over-depleting the ISM. The fraction of metals stripped as wind particles leave the galaxy in TNG is a parameter of the model, and provides a reasonable match to the observed mass-metallicity relation. This is shown in black for stellar metallicity in Fig. 17. Gas metallicity is displayed in Fig. 18, calculated in an analogous way to observations, as the ratio of oxygen to hydrogen abundances in star-forming gas cells in the simulation (without assuming oxygen makes up 35 per cent of metal mass, unlike Torrey et al., 2019).
In Arkenstone we have the opposite problem, where the low (and fixed) mass loadings tested in this paper do not remove enough metals from the ISM to match the observed mass-metallicity relation, shown in orange in Figs. 17 and 18. Both figures show that our fixed loading Arkenstone run is close to TNG at the massive end, but significantly overestimates metal masses in the dwarf regime. We note that the Arkenstone metallicity plots are will already be higher than TNG due to a larger number of stars producing more metals.
This is not unexpected, as the current implementation ejects mass with the same metallicity as the ISM; however, one might expect the hot outflows produced by supernovae to be much more enriched. Indeed, high-resolution simulations of small ISM regions find that a significant fraction of the highly enriched SN ejecta is directly launched into the wind (e.g., Hu, 2019; Kim et al., 2020a; Steinwandel et al., 2024a), such that the metallicity of the hot outflow is closer to constant rather than being directly proportional to the ISM metallicity (e.g., Li & Bryan, 2020; Kim et al., 2020b). Although we do not explore such a model in this work, we do demonstrate that the predicted mass-metallicity relation is sensitive to this assumption. In particular, we also show (in these figures) a run where we enhance the metallicity of the Arkenstone wind by a factor of 5. This lowers the normalisation of the stellar metallicities in Fig. 17, but does not notably affect the slope. The gas metallicities in Fig. 18 are lowered in low-mass galaxies while remaining largely unchanged (and consistent with observations) at the massive end. This change does not significantly affect any other galaxy properties, seen in the SMHM in Fig. 19, suggesting our other results are robust to changes in the metallicity of the wind. Note that a (nearly) constant wind metallicity would correspond to a wind metallicity enhancement factor (relative to the ISM metallicity) which increases with decreasing stellar mass, a trend that would bring our predictions into closer agreement with observations. We postpone a full study of the metal content of galaxies and the impact of varying mass loading and metal enhancements to future work, where we will simultaneously investigate the metal content of the ISM, CGM and IGM.